OpenStudy (anonymous):

I'm doing differential equations, and the problem is about oil leaking from a pipe line. My equation is dx/dt=(k+x)/x where k is a constant and x is the amount of oil. Can anyone help me out?

OpenStudy (solomonzelman):

You want to solve the DE?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{dx }{dt}=\frac{k+x}{x} }\) (so x is a function of t)

OpenStudy (solomonzelman):

You can do separation of variables.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{dx }{dt}=\frac{k+x}{x} }\) \(\color{#000000 }{ \displaystyle \frac{x}{k+x} \cdot \frac{dx }{dt}=1 }\) \(\color{#000000 }{ \displaystyle \color{#ff0000}{\int}\frac{x}{k+x} \cdot \frac{dx }{\cancel{dt}}\color{#ff0000}{\cancel{dt}}=\color{#ff0000}{\int}1 \color{#ff0000}{dt}}\)

OpenStudy (anonymous):

what I got was \[x-k \ln|x+k|+C=t\]

OpenStudy (solomonzelman):

u=k+x u-k=x integrand is: (u-k)/u du 1 - k(u)\(^{-1}\) du antideriv. u - kln(u) +C so you get; k+x - kln|x+k|+C=t and k dissolves with C. GOOD!

OpenStudy (solomonzelman):

now, you need to solve for x.

OpenStudy (anonymous):

how do i go about pulling the variables out of ln|x=k|without turning the other x into an exponent of e?

OpenStudy (solomonzelman):

you need to use the Albert .. blah (forgot the name) function W(x)

OpenStudy (solomonzelman):

It is the inverse of xe^x..

OpenStudy (solomonzelman):

this is a pretty hard equation to solve for x...

OpenStudy (solomonzelman):

that is wolfram's result. http://www.wolframalpha.com/input/?i=x-k+ln%28x%2Bk%29%2BC%3Dt++solve+for+x